OpenStudy (hy123):

Simplify (4-4w^2)/(w^2-7w+6)

OpenStudy (hy123):

\[(4(1-w^2))\div(w-1)(w-6)\]

OpenStudy (lgbasallote):

ahh so what's your problem here?

OpenStudy (hy123):

I was doublechecking to see if it was simplified enough

OpenStudy (lgbasallote):

\[\large \frac{4(1 - w^2)}{(w-1)(w-6)}\] \[\large \frac{4(1+w)(1-w)}{(w-1)(w-6)}\] \[\large \frac {-4(1+w)\cancel{(w-1)}}{\cancel{(w-1)}(w-6)}\] got it?

OpenStudy (hy123):

Got it. I knew u could factor the W^2-1

OpenStudy (lgbasallote):

take note what i did..it was originally (1+w)(1-w) but i was able to make it into (1+w)(w-1) by factoring a negative 1